# Learning from Stochastic Teacher Representations Using Student-Guided Knowledge Distillation

### v1

@@ Abstract
Advances in self-distillation have shown that when knowledge is distilled from a teacher to a student using the same deep learning (DL) architecture, the student performance can surpass the teacher particularly when the network is overparameterized and the teacher is trained with early stopping. Alternatively, ensemble learning also improves performance, although training, storing, and deploying multiple models becomes impractical as the number of models grows. Even distilling an ensemble to a single student model or weight averaging methods first requires training of multiple teacher models and does not fully leverage the inherent stochasticity for generating and distilling diversity in DL models. These constraints are particularly prohibitive in resource-constrained or latency-sensitive applications such as wearable devices. This paper proposes to train only one model and generate multiple diverse teacher representations using distillation-time dropout. However, generating these representations stochastically leads to noisy representations that are misaligned with the learned task. To overcome this problem, a novel stochastic self-distillation (SSD) training strategy is introduced for filtering and weighting teacher representation to distill from task-relevant representations only, using student-guided knowledge distillation (SGKD). The student representation at each distillation step is used as authority to guide the distillation process. Experimental results [^1] The code is made publically available at: https://github.com/haseebaslam95/SSD on real-world affective computing, wearable/biosignal datasets from the UCR Archive, the HAR dataset, and image classification datasets show that the proposed SSD method can outperform state-of-the-art methods without increasing the model size at both training and testing time, and incurs negligible computational complexity compared to state-of-the-art ensemble learning and weight averaging methods.
@@
Self Distillation Dropout Time-Series Student Guided Knowledge Distillation
@@ 1 Introduction
Self-distillation is a specialized case in KD, where the teacher and student have the same architecture, and the student typically surpasses the teacher’s performance particularly where the model is overparameterized i.e. has sufficient capacity and the teacher is trained with early-stopping. This increase in performance is typically associated with the fact that, for deep neural networks mainly, the teacher and the student have learned separate discriminative features, and self-distillation implicitly ensembles the two models [1]. Diversity in the feature space is a critical factor that enhances the robustness and accuracy of machine learning models. Diverse representations provide a comprehensive understanding of the input data, mitigating overfitting and improving generalization across various tasks [8] [15].
@@
(2) Within SSD, a novel student-guided KD (SGKD) mechanism enables the student model to selectively distill knowledge from the most informative teacher representations. Feature-based KD is used to align the student’s internal feature space with the teacher, promoting a more granular knowledge transfer.
@@ 2 Related Work
Knowledge Distillation. Originally introduced by  [4]  [11], the KD domain has evolved with several refinements in its application and architecture. Romero et al. [25] introduced the concept of distilling from feature representations instead of logits. The idea of transferring the attention maps from the teacher model to the student model was studied by Zagoruyko and Komodakis [37]. Relational KD proposed by Park et al. [23] studied the benefits of utilizing structural information for more fine-grained KD. The KD domain was extended to multi-task, semi-supervised, and unsupervised learning by Lopez et al. [19]. Yin et al. [35] proposed a data-free KD approach by generating class-conditional images.
Deep Ensembles and Model Soups. Ensembling methods improve predictive performance, generalization in neural networks, and uncertainty estimation. Deep ensemble is a simple yet effective technique where a simple aggregation of independently trained models harnesses the diversity, leading to better performance than each model. Lakshminarayanan et al. [15] demonstrated the effectiveness of deep ensembles for uncertainty estimation, showing that they outperform many Bayesian approaches in terms of both calibration and robustness. Deep theoretical insights on ensemble diversity were provided by Fort et al. [8]. Moreover, Ovadia et al. [22] highlight the advantages of deep ensembles in handling distributional shifts, reinforcing their utility in real-world scenarios. Despite their advantages, deep ensembles are computationally expensive, requiring the training and storage of multiple models, which motivates research into alternative methods that capture similar benefits with reduced complexity.
@@
Self-Distillation. This term has been used in the literature in two different contexts: distilling knowledge from deeper layers in a model to shallower layers of the same model’s instance or through the use of an auxiliary network [17, 38], and knowledge distilled from a model to another instance of the model with the same architecture [1, 9, 21]. In this work, ’self-distillation’ refers to the latter. Furlanello et al. [9] proposed born-again neural networks, a seminal work exploring KD using the same model for teacher and student, showing that the student can outperform the teacher. Iterative distillation from the trained student, used as a teacher for the subsequent student model, also improved performance. Dong et al. [7] showed that early stopping is crucial in harnessing dark knowledge in self-distillation settings. Dark knowledge is the hidden class relationships encoded in the teacher model’s soft probability outputs, which provide more information than hard labels. This nuanced information helps the student model learn better generalization and richer representations. A direct correlation between the diversity in the teacher predictions and student performance was studied in depth by Zhang et al. [39]. The authors enhanced the predictive diversity through a novel instance-specific label smoothing.
@@ 3 Proposed Method
Problem Definition: Given a trained teacher model $\mathcal{T}$, the challenge is to effectively transfer its knowledge to a student model $\mathcal{S}$ to maximize the student’s generalization performance. Standard self-distillation techniques often treat the teacher’s outputs as deterministic, failing to exploit the inherent stochasticity that can provide richer and more diverse information. On the other hand, stochastically obtaining the teacher representations introduces diversity but at the cost of generating noisy representations. The problem, therefore, is to design a KD framework that leverages the variability in the teacher’s representations, generated through stochastic mechanisms like dropout, while ensuring that the student learns task-relevant information in a computationally efficient manner. The main aim of our paper is to filter out the noisy representations from $\mathcal{F}^{\mathcal{T}}(x)=[f^{\mathcal{T}}_{1}(x),f^{\mathcal{T}}_{2}(x),\dots,f^{\mathcal{T}}_{n}(x)]$ and obtain weighted teacher representation $\hat{f}^{\mathcal{T}}(x)$ to selectively distill from the relevant teacher representations.
Figure 1: Illustration of the proposed SSD training strategy. The teacher model ($\mathcal{T}$) is trained, and its parameters are frozen except for the dropout layers in the student training stage. In SGKD, for each input $x\in\mathcal{X}$, $n$ stochastic teacher representations $f^{\mathcal{T}}_{i}(x)$ and one student representation is obtained $f^{\mathcal{S}}$, which are fed to the SSD module that outputs $\hat{f}^{\mathcal{T}}(x)$. Feature-based KD is then applied on $\hat{f}^{\mathcal{T}}(x)$ and $f^{\mathcal{S}}(x)$, with addition of $\mathcal{L}_{\text{task}}$, and the student model parameters $\theta_{\mathcal{S}}$ are updated. The part inside the dashed orange block is kept at inference.
@@ 3.2 Student-Guided Knowledge Distillation
The SSD method relies on selecting teacher representations that would mimic an ensemble of independently trained teacher models. In other words, it masks out the representations that are too far off from the current student representations. A direct way of selecting such representations would be to use the top-$k$ strategy at each distillation step. This strategy works, however, this makes k a hyperparameter and agnostic to the distribution of $f^{\mathcal{T}}_{i}(x)\in\mathcal{F}^{\mathcal{T}}$ at each distillation step. To avoid a manual selection of meaningful representations using a top-$k$ strategy, $\alpha_{i}$ is masked for all indices falling outside of the 90th percentile, denoted by $\hat{\alpha}_{i}$. This is discussed in detail in Section 4.3.
$\hat{\alpha}_{i}=\begin{cases}\alpha_{i},&\text{if }\alpha_{i}\geq\text{90th percentile},\\ 0,&\text{otherwise.}\end{cases}$ (2)
After obtaining the regularized attention weights $\hat{\alpha}_{i}$, the original teacher feature representations $f^{\mathcal{T}}_{i}(x)$ are weighed using $\hat{\alpha}_{i}$ to obtain the attended teacher feature vector $\hat{f}^{\mathcal{T}}(x)$ as:
@@
where $\lambda$ is the weighting parameter for the distillation loss. The method also allows for the additional constraint for logit-level distillation and can be added to $\mathcal{L}_{\text{total}}$. For the unsupervised/contrastive loss-based methods, we use the intermediate teacher representations to obtain the $\hat{f}^{\mathcal{T}}(x)$. In case of augmented views of the input sample, we average the distillation loss from both views.
@@ 4.1 Experiment Setup
SSD is validated on: (i) real-world affective computing datasets, the Biovid Heat Pain Database [30] and StressID dataset [5], (ii) wearable and biosignal datasets from the UCR Archive [6], (iii) Human Activity Recognition dataset set from the UCI Archive [2], and (iv) on benchmark image classification datasets. Appendix 0.A.1 provides details on these datasets, while Appendix 0.A.4 provides implementation details and results on the CIFAR-10 and CIFAR-100 datasets.
@@
UCR Archive. For the datasets in the UCR Archive, the proposed method was applied to two SOTA techniques – TS2Vec and SoftCLT – for unsupervised time-series representation learning. We follow the same experimental methodology proposed in TS2Vec [36] and SoftCLT [16]. For the TS2Vec method, the number of stochastic teacher representations was 15, with a teacher dropout rate $p$ of 0.2, and the student dropout rate was set to 0.1. The value of $H$ was 5. For loss weighting, $\lambda$ was set to 0.2. For softCLT, all the parameters were kept the same as those for TS2Vec except for the value of $H$, which was set to 15. All experiments were performed on the NVIDIA A100-SXM4-40GB GPUs.
@@ 4.2 Comparison Against State-of-the-Art Methods
Affective Computing Datasets. Table 1 reports the results of SSD against SOTA methods on Biovid. SSD improves accuracy by $2.5\%$ over the SOTA. We achieve $84.59\%$ on the Biovid dataset using the EDA modality without applying SSD, and $86.90\%$ after the proposed SSD method.
@@
SSD aims to minimize the space and computational complexity both during training and testing. Fig. 2 compares the performance gain in terms of model size at inference. We compare traditional ensembles (majority voting and averaging), stochastic weight averaging, uniform soup (uniform weight averaging), and greedy soup (a greedy approach for weight averaging). The marker size in Fig. 2 denotes the model size at inference; since traditional ensembles require storing all of the trained models for inference, it becomes impractical to deploy for inference on wearable devices. It can be observed from Fig. 2 that both majority vote and averaging-based ensembles increase the model performance, but the model size proportionally increases; essentially, for an ensemble model with 25 models, it would become 25$\times$ the size of the baseline model. On the other hand, model soups do not increase the model size at inference but are still computationally expensive in terms of train-time FLOPs as shown in Appendix 0.B. The total number of FLOPs increases from $\approx$ 0.87 G-FLOPs to $\approx$ 21.8 G-FLOPs. In contrast, the proposed method achieves comparable performance to the traditional approaches, i.e., 1.8% increase in the accuracy over the baseline, while keeping the model size the same as the baseline model, and the train-time computation is significantly less since SSD requires the model to be trained twice, once in the teacher training process and second for student training.
@@ 4.3 Ablations
Number of Stochastic Representations. As discussed before, meaningful diversity in the teacher space is crucial for the superior performance of an ensemble and, by extension, also crucial in SSD, since it also implicitly ensembles teacher representations. The number of repetitions dictates how diverse $f^{\mathcal{T}}_{i}(x)$ is. We evaluated with different settings and reported the results in Table 5. Distill All - the first two rows show the results without applying SSD and learning from all stochastic representations. This could also be seen as an alternative way of teaching students to drop out. Rows 3-6 show results with an increasing number of total repetitions and selected representations. As the total number of representations becomes too large, the performance drops even when applying SSD. This leads us to believe that when you select a more significant number of representations, the noisy representations leak through the filtering mechanism and become part of the distillation process. This problem also shows that selecting top-$k$ is not the best strategy in this case. Hence, we apply dynamic selection based on 90th percentile thresholding, which gives the best results.
@@
Dropout Rate The extent of diversity in the teacher space is directly related to the distillation-time dropout rate. To study how the probability value of each neuron to be deactivated affects the teacher representation space, we plot and compare the t-SNE plots with different dropout rates. Figs. 3(a)-(d) are plotted with dropout rates $0.1$, $0.2$, $0.5$ and $0.9$ respectively. It is observed that for smaller dropout rates (Fig. 3(a)), the teacher can maintain its discriminative ability; however, the three teacher representations $f^{\mathcal{T}}_{1}(x)$ (triangle), $f^{\mathcal{T}}_{2}(x)$ (circle), $f^{\mathcal{T}}_{3}(x)$ (square) mostly overlap each other, effectively meaning there is not enough diversity in the teacher space. For dropout rate $0.2$ (Fig. 3(b)), the three representations are diverse while maintaining the original structure, which shows that the teacher space has become diverse while maintaining the discriminative ability. In Fig. 3(c), the three reps are adequately spaced, but the model loses its discriminative ability.
@@
It can be observed from the figure that in each instance, the random initialization performs even worse than the baseline. This shows the effectiveness of the proposed method, because if the student model is not initialized with $\theta^{\mathcal{T^{\prime}}}$, the $\phi_{i}$ calculated would be ineffective since the current $f^{\mathcal{S}}(x)$ is not task aligned. Hence adding the additional constraint in the $\mathcal{L}_{total}$ term breaks the performance. See Appendix 0.C.1 for detailed results.

### v2

@@ Abstract
Advances in self-distillation have shown that when knowledge is distilled from a teacher to a student using the same deep learning (DL) model, student performance can surpass the teacher, particularly when the model is over-parameterized and the teacher is trained with early stopping. Alternatively, ensemble learning also improves performance, although training, storing, and deploying multiple DL models becomes impractical as the number of models grows. Even distilling a deep ensemble to a single student model or weight averaging methods first requires training of multiple teacher models and does not fully leverage the inherent stochasticity for generating and distilling diversity in DL models. These constraints are particularly prohibitive in resource-constrained or latency-sensitive applications on, e.g., wearable devices. This paper proposes to train only one model and generate multiple diverse teacher representations using distillation-time dropout. However, generating these representations stochastically leads to noisy representations that are misaligned with the learned task. To overcome this problem, a novel stochastic self-distillation (SSD) training strategy is introduced for filtering and weighting teacher representation to distill from task-relevant representations only, using student-guided knowledge distillation. The student representation at each distillation step is used to guide the distillation process. Experimental results [^1] Code and supplementary available at: https://github.com/haseebaslam95/SSD on real-world affective computing, wearable/biosignal (UCR Archive), HAR, and image classification datasets show that the proposed SSD method can outperform state-of-the-art methods without increasing the model size at both training and testing time. It incurs negligible computational complexity compared to ensemble learning and weight averaging methods.
@@
Deep Learning Self Distillation Dropout Time-Series Student-Guided Knowledge Distillation
@@ 1 Introduction
Self-distillation is a specialized case in KD, where the teacher and student have the same DL architecture, and the student typically surpasses the teacher’s performance particularly where the model is over-parameterized i.e., has sufficient capacity and the teacher is trained with early-stopping. This increase in performance is typically associated with the fact that, with DL models, the teacher and the student have learned separate discriminative features, and self-distillation implicitly ensembles the two models [1]. Diversity in the feature space is a critical factor that enhances the robustness and accuracy of machine learning models. Diverse representations provide a comprehensive understanding of the input data, mitigating overfitting and improving generalization across various tasks [9] [17].
@@
(2) Within SSD, a novel SGKD mechanism enables the student model to selectively distill knowledge from the most informative teacher representations. Feature-based KD is used to align the student’s internal feature space with the teacher, promoting a more granular knowledge transfer.
@@ 2 Related Work
Knowledge Distillation. Originally introduced by  [5]  [12], the KD domain has evolved with several refinements in its application and architecture. Romero et al. [29] introduced the concept of distilling from feature representations instead of logits. The idea of transferring the attention maps from the teacher model to the student model was studied by Zagoruyko and Komodakis [41]. Relational KD proposed by Park et al. [27] studied the benefits of utilizing structural information for more fine-grained KD. The KD domain was extended to multi-task, semi-supervised, and unsupervised learning by Lopez et al. [23]. KD has also been studied in multimodal systems, particularly with applications like cross-modal KD [30] privileged KD [3], federated learning [21], are a few examples of the widespread application of KD in real-world systems.
Deep Ensembles and Model Soups. Ensembling methods improve predictive performance, generalization in neural networks, and uncertainty estimation. Deep ensemble is a simple yet effective technique where a simple aggregation of independently trained models harnesses the diversity, leading to better performance than each model. Lakshminarayanan et al. [17] demonstrated the effectiveness of deep ensembles for uncertainty estimation, showing that they outperform many Bayesian approaches in terms of both calibration and robustness. Deep theoretical insights on ensemble diversity were provided by Fort et al. [9]. Moreover, Ovadia et al. [26] highlight the advantages of deep ensembles in handling distributional shifts, reinforcing their utility in real-world scenarios. Despite their advantages, deep ensembles are computationally expensive, requiring the training and storage of multiple models, which motivates research into alternative methods that capture similar benefits with reduced complexity. More recently, parameter-efficient fine-tuning techniques like low-rank adaptation (LoRA) [13] have enabled efficient fine-tuning of large models. For example, Li et al. [20] introduced Ensembles of Low-Rank Expert Adapters. However, these techniques still require i) careful adaptation of each model, and ii) storing all the models for inference.
@@
Self-Distillation. This term has been used in the literature in two different contexts: distilling knowledge from deeper layers in a model to shallower layers of the same model’s instance or through the use of an auxiliary network [19, 42], and knowledge distilled from a model to another instance of the model with the same architecture [1, 10, 25]. In this work, ’self-distillation’ refers to the latter. Furlanello et al. [10] proposed born-again neural networks, a seminal work exploring KD using the same model for teacher and student, showing that the student can outperform the teacher. Iterative distillation from the trained student, used as a teacher for the subsequent student model, also improved performance. Dong et al. [dong2019] showed that early stopping is crucial in harnessing dark knowledge in self-distillation settings. Dark knowledge is the hidden class relationships encoded in the teacher model’s soft probability outputs, which provide more information than hard labels. This nuanced information helps the student model learn better generalization and richer representations. A direct correlation between the diversity in the teacher predictions and student performance was studied in depth by Zhang et al. [43]. The authors enhanced the predictive diversity through a novel instance-specific label smoothing.
@@ 3 Proposed Method
Problem Definition: Given a trained teacher model $\mathcal{T}$, the challenge is to effectively transfer its knowledge to a student model $\mathcal{S}$ such that its generalization performance is maximized. Standard self-distillation techniques often treat the teacher’s outputs as deterministic, failing to exploit the inherent stochasticity that can provide richer and more diverse information. On the other hand, stochastically obtaining the teacher representations introduces diversity but at the cost of generating noisy representations. The problem, therefore, is to design a KD framework that leverages the variability in the teacher’s representations, generated through stochastic mechanisms like dropout, while ensuring that the student learns task-relevant information in a computationally efficient manner. The main aim of our paper is to filter out the noisy representations from $\mathcal{F}^{\mathcal{T}}(x)$ and obtain weighted teacher representation $\hat{f}^{\mathcal{T}}(x)$ to selectively distill from the relevant teacher representations.
Figure 1: Illustration of the proposed SSD training strategy. The teacher $\mathcal{T}$ is trained, and its parameters are frozen except for the dropout layers in the student training stage. In SGKD, for each input $x\in\mathcal{X}$, $n$ stochastic teacher representations $f^{\mathcal{T}}_{i}(x)$, $i=1,2,...,n$, and one student representation is obtained $f^{\mathcal{S}}$, which are fed to the SSD module that outputs $\hat{f}^{\mathcal{T}}(x)$. Feature-based KD is then applied on $\hat{f}^{\mathcal{T}}(x)$ and $f^{\mathcal{S}}(x)$, with addition of $\mathcal{L}_{\text{task}}$, and the student model parameters $\theta_{\mathcal{S}}$ are updated. The part inside the dashed orange block is kept at inference.
@@ 3.2 Student-Guided Knowledge Distillation
The SSD method relies on selecting teacher representations that would mimic an ensemble of independently trained teacher models. In other words, it masks out the representations that are too different from the current student representations. A direct way of selecting such representations would be to use the top-$k$ strategy at each distillation step. Although this strategy can work, it relies on k a hyperparameter that is agnostic to the distribution of $f^{\mathcal{T}}_{i}(x)\in\mathcal{F}^{\mathcal{T}}$ at each distillation step. To avoid a manual selection of meaningful representations using a top-$k$ strategy, $\alpha_{i}$ is masked for all indices falling outside of the $\epsilon$-th percentile, denoted by $\hat{\alpha}_{i}$.
$\hat{\alpha}_{i}=\begin{cases}\alpha_{i},&\text{if }\alpha_{i}\geq\text{$\epsilon$},\\ 0,&\text{otherwise.}\end{cases}$ (2)
where $\epsilon\in[0,100]$ is the threshold value for masking. Section 4.3 provides a more detailed discussion. After obtaining the regularized attention weights $\hat{\alpha}_{i}$, the original teacher feature representations $f^{\mathcal{T}}_{i}(x)$ are weighed using $\hat{\alpha}_{i}$ to obtain the attended teacher feature vector $\hat{f}^{\mathcal{T}}(x)$ as:
@@
where $\lambda$ is the weighting parameter for the distillation loss. The method also allows for the additional constraint for logit-level distillation and can be added to $\mathcal{L}_{\text{total}}$. For the unsupervised/contrastive loss-based methods, we use the intermediate teacher representations to obtain the $\hat{f}^{\mathcal{T}}(x)$. In case of augmented views of the input sample, we average the distillation loss from both views. The pseudo-code for the SSD training procedure is shown in Algorithm 1.
Algorithm 1 - SSD Training Procedure.
2:Trained student model $\mathcal{S}$.
5:for each input sample $x\in\mathcal{X}$ do
9:     Mask $\alpha_{i}$ for all indices outside the $\epsilon$-th percentile using Eq. (2)
11:     Compute distillation loss $\mathcal{L}_{\text{dist}}$ using Eq. (4)
12:end for
13:Compute total loss $\mathcal{L}_{\text{total}}$ using Eq. (5)
14:Update parameters $\theta^{\mathcal{S}}$ using backpropagation
@@ 4.1 Experiment Setup
SSD is validated on: (i) real-world affective computing datasets: the Biovid Heat Pain Database [3] and StressID dataset [4], (ii) wearable and biosignal datasets from the UCR Archive [1], (iii) Human Activity Recognition (HAR) dataset set from the UCI Archive [2], and (iv) on benchmark image classification datasets. Appendix A.1 provides details on these datasets, while Appendix A.3 provides implementation details and results on the CIFAR-10 and CIFAR-100 datasets.
@@
UCR Archive. For the datasets in the UCR Archive, the proposed method was applied to two SOTA techniques – TS2Vec and SoftCLT – for unsupervised time-series representation learning. We follow the same experimental methodology proposed in TS2Vec [40] and SoftCLT [18]. For the TS2Vec method, the number of stochastic teacher representations was 15, with a teacher dropout rate $p$ of 0.2, and the student dropout rate was set to 0.1. The value of $H$ was 5. For loss weighting, $\lambda$ was set to 0.2. For softCLT, all the parameters were kept the same as those for TS2Vec except for the value of $H$, which was set to 15.
Computing infrastructure. All experiments were performed on the NVIDIA A100-SXM4-40GB GPUs with the $\epsilon$ value of 90.
@@ 4.2 Comparison Against State-of-the-Art Methods
Affective Computing Datasets. Table 1 reports the results of SSD and SOTA methods on Biovid. SSD improves accuracy by $2.5\%$ over the selected baseline (PAN without SSD) and $1.7\%$ over the current SOTA. Specifically, accuracy on Biovid increases from $84.59\%$ (using the EDA modality without SSD) to $86.90\%$ with the proposed SSD method.
@@
SSD aims to minimize the space and computational complexity both during training and testing. Fig. 2 compares the performance gain in terms of model size at inference. We compare traditional ensembles (majority voting and averaging), stochastic weight averaging, uniform soup (uniform weight averaging), and greedy soup (a greedy approach for weight averaging). The marker size in Fig. 2 denotes the model size at inference; since traditional ensembles require storing all of the trained models for inference, it becomes impractical to deploy for inference on wearable devices. It can be observed from Fig. 2 that both majority vote and averaging-based ensembles increase the model performance, but the model size proportionally increases; essentially, for an ensemble model with 25 models, it would become 25$\times$ the size of the baseline model. On the other hand, model soups do not increase the model size at inference but are still computationally expensive in terms of train-time FLOPs as shown in Appendix B. The total number of FLOPs increases from $\approx$ 0.87 G-FLOPs to $\approx$ 21.8 G-FLOPs. In contrast, the proposed method achieves comparable performance to the traditional approaches, i.e., 1.8% increase in the accuracy over the baseline, while keeping the model size the same as the baseline model, and the train-time computation is significantly less since SSD requires the model to be trained twice, once in the teacher training process and second for student training.
@@ 4.3 Ablations
Number of Stochastic Representations. As discussed before, meaningful diversity in the teacher space is crucial for the superior performance of an ensemble and, by extension, also crucial in SSD, since it also implicitly ensembles teacher representations. The number of repetitions dictates how diverse $f^{\mathcal{T}}_{i}(x)$ is. We evaluated with different settings and reported the results in Table 5. Distill All - the first two rows show the results without applying SSD and learning from all stochastic representations. This could also be seen as an alternative way of teaching students to drop out. Rows 3-6 show results with an increasing number of total repetitions and selected representations. As the total number of representations becomes too large, the performance drops even when applying SSD. This leads us to believe that when you select a more significant number of representations, the noisy representations bypass through the filtering mechanism and become part of the distillation process. This problem also indicates a simple top-$k$ selection is not the best strategy in this case. Hence, dynamic selection is applied based on $\epsilon$-th percentile thresholding.
@@
Dropout Rate The extent of diversity in the teacher space is directly related to the distillation-time dropout rate. To study how the probability value of each neuron to be deactivated affects the teacher representation space, we plot and compare the t-SNE plots with different dropout rates. Figs. 3(a)-(d) are plotted with dropout rates $0.1$, $0.2$, $0.5$ and $0.9$ respectively. It is observed that for smaller dropout rates (Fig. 3(a)), the teacher can maintain its discriminative ability; however, the three teacher representations $f^{\mathcal{T}}_{1}(x)$ (triangle), $f^{\mathcal{T}}_{2}(x)$ (circle), $f^{\mathcal{T}}_{3}(x)$ (square) mostly overlap each other, effectively meaning there is not enough diversity in the teacher space. For dropout rate $0.2$ (Fig. 3(b)), the three representations are diverse while maintaining the original structure, which shows that the teacher space has become diverse while maintaining the discriminative ability. In Fig. 3(c), the three representations are adequately spaced, but the model loses its discriminative ability.
@@
To further analyze the impact dropout rate on both variance and overall performance, we conducted an ablation study on the HAR dataset. Fig. 4 shows the effect of dropout rates on the student performance (dashed/black line) and the variance across teacher representations (colored/solid lines). Variance increases with the teacher dropout rate. For each dropout rate, the highest variance is observed for the lowest number of repetitions. Conversely, the lowest variance for each dropout rate is observed with the highest number of reps. Results suggest that when the number of forward passes is greater, the overall variance decreases. It may also indicate a more accurate approximation of the true variance because it is calculated with more samples. In the latter case, it can be observed that for lower dropout rates, the variances across different number of repetitions are more accurate approximation of the true variance. In terms of performance, the model peaks at a dropout rate of 0.2, and the performance starts deteriorating beyond a dropout rate of 0.5. This phenomenon can be further explained from Fig. 3 where the t-SNE plots show the model starting to lose its discriminative ability at higher dropout rates.
Figure 4: Effect of various dropout rates on the performance and variance quantification among teacher representations. Accuracy (black/dashed line), and variance for various no. of reps (colored/solid lines) are plotted against dropout rates on the HAR dataset
@@
It can be observed from the figure that in each instance, the random initialization performs even worse than the baseline. This shows the effectiveness of the proposed method, because if the student model is not initialized with $\theta^{\mathcal{T^{\prime}}}$, the $\phi_{i}$ calculated would be ineffective since the current $f^{\mathcal{S}}(x)$ is not task aligned. Hence adding the additional constraint in the $\mathcal{L}_{total}$ term breaks the performance. See Appendix C.1 for detailed results.